OpenStudy (anonymous):

If you burn 42.5 g of hydrogen and produce 3.80 × 102 g of water, how much oxygen reacted?

OpenStudy (alphadxg):

Hydrogen+Oxygen->Water If you burn 39.7 g of hydrogen and produce 355g of water, how much oxygen reacted? so you have 355 grams of water in the end or the process Find the g/mol of H2O 355g H2O *(Oxygen g/mol/H2O g/mol) = grams of oxygen you reacted. To find the g/mol of H2O simply (hydrogen g/mol)*2 + (g/mol of oxygen) = H2O g/mol 355g H2O *(15.9994 g/mol/18.0153 g/mol) = 315.28 grams of Oxygen were reacted theoretically of course assuming all the oxygen was pushed forward to water. if you look at the formula you will notice the g/mol cross out and the H2O and the H2O cross out so all you are left with is Oxygen and grams

OpenStudy (anonymous):

thank you